OpenStudy (anonymous):

find x and y, EXPLAIN :D : 3x-2y=14 2x+3y=18

OpenStudy (luffingsails):

Love the name... You can do this two ways depending on your level. One way is to solve one of the equations for one of the variables (x or y). Then substitute and solve in the other. Let's try that. Start with: 3x-2y=14 and solve for x. 3x-2y=14 3x = 14+2y x = (14+2y)/3 Now, we'll substitute that into our other equation for x. 2x + 3y = 18 2((14+2y)/3) + 3y = 18 (28 + 4y)/3 + 3y = 18 (28+4y)/3 + 9y/3 = 18 (getting a common denominator) (28 + 4y + 9y)/3 = 18 28 + 13y = 54 (after simplification) 13y = 26 y = 2. Now, take y = 2 and put it into either equation to solve for x. 2x + 3(2) = 18 2x + 6 = 18 2x = 12 x = 6. So, x = 6, y = 2. All done.